OpenStudy (anonymous):

Will someone help me figure this out? The charges on two metallic balls are 5.0 and 7.0 coulombs respectively. They are kept 1.2 meters apart. What is the force of interaction between the two? (k = 9.0 * 10^9 newton·meter^2/coulomb^2) A) 2.2 x 10^11 newtons B) 4.1 x 10^-7 newtons C) 5.0 x 10^11 newtons D) 7.0 x 10^-7 newtons

OpenStudy (anonymous):

Have you learned Coulomb's law?

OpenStudy (anonymous):

No not yet. I saw the formula but I don't get what it means o-o

OpenStudy (anonymous):

F =k q1* q2/r^2 This is it right?

OpenStudy (anonymous):

Cool cool - so Coulomb's law is \[ |\textbf F|=k\frac{q_1q_2}{r^2}\] which gives the magnitude of the force between two charged objects q1 = charge of first object q2 = charge of second object r^2 = square of the distance between them k = Coulomb's constant (whose value they give)

OpenStudy (anonymous):

yeah, what you wrote is right ^_^

OpenStudy (anonymous):

So q1=5.0 and q2=7.0 Multiply that right? And I would get 35.0 then I would divide that by 1.2 and that's 29.1666666667

OpenStudy (anonymous):

Then what o_o I dont get what I do after

OpenStudy (anonymous):

(1.2m)^2 = 1.44m^2

OpenStudy (anonymous):

Then you multiply 35C^2/1.44m^2 by k k = 9x10^9 C^2/m^2

OpenStudy (anonymous):

What.. Now youre confusing me ._. So I didnt do that right?

OpenStudy (anonymous):

not quite - remember it's \[F=k\frac{q_1q_2}{r^2}\] you got the \[q1*q2 = 7C*5C = 35C^2\] but you divide by r^2, not just r \[r^2 = (1.2m)^2 = 1.44m^2\]

OpenStudy (anonymous):

But where does the C^2 come in?

OpenStudy (anonymous):

Wait, what the heck is C? O.o

OpenStudy (anonymous):

That's the units on the charges - 5 Coulombs = 5C 7 Coulombs = 7C and when you multiply units, they acts just like normal parts of the equation so \[1C * 1C = 1C^2\] likewise \[5C*7C=3C^2\]

OpenStudy (anonymous):

But 5*7 is 35..

OpenStudy (anonymous):

Im sorry Im not really catching it .-.

OpenStudy (anonymous):

Sorry, mistype - 35C^2 you're right

OpenStudy (anonymous):

So like, as an example when you're finding the area of a rectangle  What would the answer for the area be?

OpenStudy (anonymous):

(this rectangle is in no way related to anything in this problem - it's just an example to look at multiplying units)

OpenStudy (anonymous):

4x6 = 24 o

OpenStudy (anonymous):

yah, but the units on it are \[A=(4m)(6m) = 24m^2\] right?

OpenStudy (anonymous):

Oh oh okay Im getting this xD

OpenStudy (anonymous):

^_^ !

OpenStudy (anonymous):

9x10^9 C^2/m^2 So that^ I would just do that equation then? And thats my answer o-o

OpenStudy (anonymous):

Coulomb's law is what gives the answer you're looking for \[F=k\frac{q_1q_2}{r^2}=(9.0\times 10^9 N m^2/C^2)\frac{35C^2}{1.44m^2}\] See how in the units, C^2 cancels, m^2 cancels, so you're only left with Newtons for force?

OpenStudy (anonymous):

35/1.44?

OpenStudy (anonymous):

24.30??

OpenStudy (anonymous):

I should just divide those to numbers then? or those numbers cancel each other out too?

OpenStudy (anonymous):

You divide the numbers, the units are what cancels out

OpenStudy (anonymous):

\[F=\left((9\times 10^9)\frac{35}{1.44}\right)\left((N\cancel{m^2}/\cancel{C^2})\frac{\cancel{C^2}}{\cancel{m^2}}\right)\] \[=\big(2.19\times10^{11}\big)\big(N\big)\]

OpenStudy (anonymous):

Oh. So all I had to do was that ._. Atleast i know how to do it know for next time. Thank you very much.

OpenStudy (anonymous):

Sorry to complicate things :/

OpenStudy (anonymous):

Oh no no. You helped me a bunch [: You have no idea xD